# FedTLU: Federated Learning with Targeted Layer Updates

### v1

@@ Abstract
Federated learning (FL) addresses privacy concerns in language modeling by enabling multiple clients to contribute to training language models. However, non-IID (identically and independently distributed) data across clients often limits FL’s performance. This issue is especially challenging during model fine-tuning, as noise due to variations in clients’ data distributions can harm model convergence near the optimum. This paper proposes a targeted layer update strategy for fine-tuning in FL. Instead of randomly updating layers of the language model, as often done in practice, we use a scoring mechanism to identify and update the most critical layers, avoiding excessively noisy or even poisoned updates by freezing the parameters in other layers. We show in extensive experiments that our method improves convergence and performance in non-IID settings, offering a more efficient approach to fine-tuning federated language models.
@@ I Introduction
Language modeling has driven major application advancements in natural language processing like conversational agents and machine translation [1, 2, 3, 4]. However, the widespread use of language data raises privacy concerns, as language data from individuals is personal and prone to exploitation for malicious purposes[5, 6, 4].
@@
Federated Learning (FL) addresses these concerns by training a global model without centralizing user data [7, 8, 9]. However, one of the core challenges in FL is developing a global model that performs well across clients, whose data can vary significantly in size and distribution [10, 11]. This non-IID nature of client data often causes federated models to underperform compared to centralized models [12, 13, 10]. Language modeling has these issues because clients may exhibit distinct linguistic patterns, content, and topics, further complicating model convergence and generalization [14, 15, 16]. Toward the end of the training, as the model nears a stationary point, stochastic local updates—amplified by data heterogeneity—create variability that makes it difficult to maintain global model consistency [17, 18]. This variability slows convergence and prevents the global model from reaching optimal performance [17, 18, 19, 20].
@@
While local fine-tuning helps models adapt to client-specific data [21, 22], fine-tuning only local models can lead to divergence, as each client’s model becomes too specialized to its local data and tends to overfit, reducing the global model’s overall effectiveness across all clients [23]. Moreover, individual clients often lack enough data for effective independent fine-tuning [24, 25]. A well-optimized global model, updated across clients, provides a stronger starting point [26]. Once aligned, local fine-tuning can enhance specific client tasks without diverging from the global objective. Therefore, FL requires a well-generalized global model to meet the collective needs of all participants. Therefore, the primary focus should be on fine-tuning the global model to ensure consistency across clients, with local tuning as an optional step for personalization.
This work aims to improve global model performance in FL. Uniformly updating all layers during aggregation, especially in the final stages, can amplify noise from certain clients, degrading performance and hindering convergence [27, 28]. Toward the end of training, as the model nears a stationary point and gradients shrink, noisy or non-representative updates from certain clients can disproportionately affect the global model [8].
@@
Therefore, we can mitigate the impact of such noisy updates by selectively updating only the most important parameter sets, such as layers, which are likely to significantly reduce global loss. However, the key challenge in this approach is determining which layers to update. Clients do not know the updates made by other clients and have limited information about the criticality of each layer. Therefore, server-side coordination is needed to aggregate client information and determine the most critical layers to update while considering the heterogeneous data distributions across clients.
@@
$\bullet$ FedTLU consistently outperforms random and last-layer updates, achieving better test performance up to 7.86% globally and up to 8.27% locally, particularly during the noisy fine-tuning phase.
We review related work in the ‘Related Work’ section, present the FedTLU algorithm in ‘Methodology,’ and offer theoretical support in ‘Theoretical Analysis.’ The ‘Experimental Evaluation’ validates our approach, and we summarize our findings in the ‘Conclusion.’
@@ III Methodology
This section introduces FedTLU, a method for selecting layers to update during global aggregation in FL, particularly for fine-tuning language models with non-IID data. The goal is to efficiently update the most critical layers or blocks to improve model convergence and generalization, especially in the final training stages.
@@
where $\|\Delta W_{i}\|$ is the $L2$ norm of the difference between the parameters before round $t$’s update $(W_{i}^{(t)})$ and the aggregated parameters after the update $W_{\text{aggre},i}^{(t+1)}$, $n_{i}$ is the number of parameters in layer $W_{i}$, $\text{std}(\Delta W_{i})$ is the standard deviation of the differences in parameters in the layer.
@@ IV-A Assumptions
The gradients of the global loss function are Lipschitz continuous, implying that the global loss function is smooth. Specifically, for the aggregated global model $W$, the loss function $\mathcal{L}(W)$ is $L$-smooth, meaning:
@@
We assume that the subset of layers $W_{\mathcal{S}}^{(t)}$ contains layers where the gradient most effectively reduces the loss. Specifically, we assume:
@@
For simplicity, we assume full participation from all clients in every round of communication. Additionally, we assume that each client’s local data distribution remains constant throughout the training process.
@@ IV-B Analysis of Gradient Descent Updates
When the entire model $W$ is updated during the global aggregation phase, the update rule for each layer $W_{i}$ at iteration $t$ is
$W_{i}^{(t+1)}=W_{i}^{(t)}-\eta\nabla_{W_{i}}\mathcal{L}(W^{(t)}),$
where $\nabla_{W_{i}}\mathcal{L}(W^{(t)})=\sum_{k=1}^{K}p_{k}\nabla_{W_{i}}\mathcal{L}_{k}(W^{(t)})$. The overall change in the global loss function due to this update is governed by the smoothness condition, which can be expressed as:
@@
where $\eta$ is the learning rate, $\|\nabla\mathcal{L}(W^{(t)})\|^{2}$ represents the squared $L2$ norm of the global gradient $\nabla\mathcal{L}(W)^{(t)}$, $L$ is the Lipschitz constant of the global loss function.
@@
Thus, the full model can be expressed as $W^{(t)}=(W^{(t)}_{\mathcal{S}},W^{(t)}_{\mathcal{S}^{c}})$, where $\mathcal{S}^{c}$ denotes the complement of $\mathcal{S}$. Since the update only affects $W_{\mathcal{S}}$, the difference $W^{(t+1)}-W^{(t)}$ simplifies to : $(W_{\mathcal{S}}^{(t+1)}-W_{\mathcal{S}}^{(t)},0)$. Substituting this into the smoothness condition:
@@
Now, let’s focus on the gradient $\nabla\mathcal{L}(W^{(t)})$. This gradient is with respect to the entire model $W^{(t)}$, which means it includes components for both $W_{\mathcal{S}}^{(t)}$ and $W_{\mathcal{S}^{c}}^{(t)}$:
$\langle\nabla\mathcal{L}(W^{(t)}),(W_{\mathcal{S}}^{(t+1)}-W_{\mathcal{S}}^{(t)},0)\rangle=\langle\nabla_{W_{\mathcal{S}}^{(t)}}\mathcal{L}(W^{(t)}),W_{\mathcal{S}}^{(t+1)}-W_{\mathcal{S}}^{(t)}\rangle.$
@@
Assuming that the gradient over the subset $W_{\mathcal{S}}^{(t)}$ is not aligned with the full gradient, i.e., large $\delta$. The loss reduction is defined as:
@@
For the subset update can be expressed as:
@@ V-A Experimental Setup and Evaluation Metrics
The network architectures consist of the repeated (main body) and non-repeated parts (input and output layers). Four update strategies were evaluated: Full, FedTLU, Random, and Last updates. In the Full update, all layers were updated in each round. For Random and FedTLU, updates were applied to the non-repeated part and a portion of the main body layers, with FedTLU selecting layers based on scores and Random selecting layers randomly. The Last update case involved updating only the last few layers of the non-repeated part.
@@ V-B Experimental Results
For training from scratch, each communication round varied the portion of updated layers (75%, 50%, and 25%). Each local client was allocated up to a maximum number of tokens, calculated by dividing the total dataset size by the number of clients. A minimum number of tokens is half of this maximum number. Local datasets were non-overlapping, and 10% of clients participated per round. Table I shows that FedTLU consistently achieved better test perplexity, reducing global perplexity by up to 1.14% and local perplexity by up to 1.3% compared to Random. FedTLU also showed a 41.7-120.8% improvement in global perplexity and 13.8-116.2% in local perplexity compared to last-layer updates. These results indicate that targeted layer updates in FedTLU enhance model stability.
For fine-tuning, we trained the global model for 400 rounds with full updates. During fine-tuning, 5% of clients participated per round, and each client’s local data was halved. The update portion was decreased sequentially from 75% to 50% to 25%, and if the loss did not decrease by at least 1% for 10 consecutive rounds, the portion was further reduced. Table II and Fig. 1 show that FedTLU outperformed other strategies, reducing global perplexity by 2.54-3.50% compared to full updates, by up to 0.97% compared to Random, and by 3.18-3.80% compared to last-layer updates. Locally, FedTLU achieved reductions of 2.60-3.15%, 0.06-0.92%, and 3.23-3.36%. In Fig. 1, FedTLU consistently maintained lower global perplexity during training.
@@ VI Conclusion
VII Acknowledgment
The authors are partially supported by the National Science Foundation under grants CNS-2409138 and CNS-2106891 and by the Department of Energy under grant DESC0025652.

### v2

# Pittsburgh, USA

@@ Abstract
Federated learning (FL) addresses privacy concerns in training language models by enabling multiple clients to contribute to the training, without sending their data to others. However, non-IID (identically and independently distributed) data across clients often limits FL’s performance. This issue is especially challenging during model fine-tuning, as noise due to variations in clients’ data distributions can harm model convergence near stationary points. This paper proposes a targeted layer update strategy for fine-tuning in FL. Instead of randomly updating layers of the language model, as often done in practice, we use a scoring mechanism to identify and update the most critical layers, avoiding excessively noisy or even poisoned updates by freezing the parameters in other layers. We show in extensive experiments that our method improves convergence and performance in non-IID settings, offering a more efficient approach to fine-tuning federated language models.
@@ I Introduction
Data-driven language models have recently driven major application advancements in natural language processing like conversational agents and machine translation [1, 2, 3, 4]. However, the widespread use of language data to train these models raises privacy concerns, as language data from individuals is personal and prone to exploitation for malicious purposes[5, 6, 4].
@@
Federated Learning (FL) addresses these concerns by training a global model without centralizing user data [7, 8, 9]. However, one of the core challenges in FL is developing a global model that performs well across clients, whose data can vary significantly in size and distribution [10, 11]. This non-IID nature of client data often causes federated models to underperform compared to centralized models [12, 13, 10]. In language models, this heterogeneity can manifest as clients’ exhibiting distinct linguistic patterns, content, and topics, complicating model convergence and generalization [14, 15, 16]. Toward the end of the training, as the model nears a stationary point, stochastic local updates—amplified by data heterogeneity—create variability that makes it difficult to maintain global model consistency [17, 18]. This variability slows convergence and prevents the global model from reaching optimal performance [17, 18, 19, 20].
@@
While local fine-tuning helps models adapt to client-specific data [21, 22], fine-tuning only local models can lead to divergence, as each client’s model becomes too specialized to its local data and tends to overfit, reducing the global model’s overall effectiveness across all clients [23]. Moreover, individual clients often lack enough data for effective independent fine-tuning [24, 25]. A well-optimized global model, updated across clients, provides a stronger starting point [26]. Once aligned, local fine-tuning can enhance specific client tasks without diverging from the global objective. Therefore, FL requires a well-generalized global model to meet the collective needs of all participants. Thus, our primary focus in this work is on fine-tuning the global model to ensure consistency across clients, with local tuning as an optional step for personalization.
@@
In FL, a central server periodically aggregates locally updated client models. However, uniformly updating all model layers during aggregation, especially in the final stages, can amplify noise from certain clients, degrading performance and hindering convergence [27, 28]. Toward the end of training, as the model nears a stationary point and gradients shrink, noisy or non-representative updates from certain clients can disproportionately affect the global model [8]. We can mitigate the impact of such noisy updates by selectively updating only the most important parameter sets, such as layers, which are likely to significantly reduce global loss. However, the key challenge in this approach is determining which layers to update. Clients do not know the updates made by other clients and have limited information about the criticality of each layer. Therefore, server-side coordination is needed to aggregate client information and determine the most critical layers to update, though the server has only limited knowledge of the heterogeneous data distributions across clients.
@@
$\bullet$ FedTLU consistently outperforms random and last-layer updates, achieving better test performance by up to 7.86% globally and up to 8.27% locally, particularly during the noisy fine-tuning phase.
We review related work in Section II, present the FedTLU algorithm in Section III, and offer theoretical support for FedTLU’s design in Section IV. Section V presents experimental results that validate our approach, and we summarize our findings in Section VI.
@@ III Methodology
This section introduces FedTLU, a method for selecting layers to update during global aggregation in FL, particularly for training language models with non-IID data. The goal is to efficiently update the most critical layers or blocks to improve model convergence and generalization, especially in the final training stages.
@@
where $\|\Delta W_{i}\|$ is the $L2$ norm of the difference between the parameters before round $t$’s update $(W_{i}^{(t)})$ and the aggregated parameters after the update $W_{\text{aggre},i}^{(t+1)}$, $n_{i}$ is the number of parameters in layer $W_{i}$, and $\text{std}(\Delta W_{i})$ is the standard deviation of the differences in parameters in layer $i$. This score is large in two cases:
@@ IV-A Assumptions
The gradients of the global loss function are Lipschitz continuous, implying that the global loss function is smooth. Specifically, for any aggregated global model $W$, the loss function $\mathcal{L}(W)$ is $L$-smooth, meaning:
@@
We assume that the subset of layers $W_{\mathcal{S}}^{(t)}$ contains layers where the gradient most effectively reduces the loss. Specifically, we assume that in each round $t$:
@@
For simplicity, we assume full participation from all clients in every round of communication. To remove stochasticity from the analysis, we also assume that each client’s local update is computed with one step of gradient descent with the full gradient of its local loss over all of its data in every round. Additionally, we assume that each client’s local data distribution remains constant throughout the training process.
@@ IV-B Analysis of Gradient Descent Updates
When the entire model $W$ is updated during the global aggregation phase, the update rule for each layer $W_{i}$ at iteration $t$ is $W_{i}^{(t+1)}=W_{i}^{(t)}-\eta\nabla_{W_{i}}\mathcal{L}(W^{(t)}),$ where $\nabla_{W_{i}}\mathcal{L}(W^{(t)})=\sum_{k=1}^{K}p_{k}\nabla_{W_{i}}\mathcal{L}_{k}(W^{(t)})$. The overall change in the global loss function due to this update is governed by the smoothness condition, which can be expressed as:
@@
where $\eta$ is the learning rate, $\|\nabla\mathcal{L}(W^{(t)})\|^{2}$ represents the squared $L2$ norm of the global gradient $\nabla\mathcal{L}(W)^{(t)}$, and $L$ is the Lipschitz constant of the global loss function.
@@
Thus, without loss of generality the full model can be expressed as $W^{(t)}=(W^{(t)}_{\mathcal{S}},W^{(t)}_{\mathcal{S}^{c}})$, where $\mathcal{S}^{c}$ denotes the complement of $\mathcal{S}$. Since the update only affects $W_{\mathcal{S}}$, the difference $W^{(t+1)}-W^{(t)}$ simplifies to : $(W_{\mathcal{S}}^{(t+1)}-W_{\mathcal{S}}^{(t)},0)$. Substituting into the smoothness condition:
@@
Now, let’s focus on the gradient $\nabla\mathcal{L}(W^{(t)})$. This gradient is with respect to the entire model $W^{(t)}$, which means it includes components for both $W_{\mathcal{S}}^{(t)}$ and $W_{\mathcal{S}^{c}}^{(t)}$, leading to $\langle\nabla\mathcal{L}(W^{(t)}),(W_{\mathcal{S}}^{(t+1)}-W_{\mathcal{S}}^{(t)},0)\rangle=\langle\nabla_{W_{\mathcal{S}}^{(t)}}\mathcal{L}(W^{(t)}),W_{\mathcal{S}}^{(t+1)}-W_{\mathcal{S}}^{(t)}\rangle$.
@@
Let us assume that the gradient over the subset $W_{\mathcal{S}}^{(t)}$ is not aligned with the full gradient, i.e., large $\delta$. The loss reduction is defined as:
@@
For the subset update, the loss reduction can be expressed as:
@@ V-A Experimental Setup and Evaluation Metrics
The network architectures consist of the repeated (main body) and non-repeated parts (input and output layers). Four update strategies were evaluated: Full, FedTLU, Random, and Last updates. In the Full update, all layers were updated in each round. For Random and FedTLU, updates were applied to the non-repeated part and a portion of the main body layers, with FedTLU selecting layers based on scores and Random selecting layers uniformly at random. The Last update case involved updating only the last few layers of the non-repeated part. Importantly, for FedTLU, Random, and Last updates, the same number of layers were updated in each experiment to ensure a fair comparison.
@@ V-B Experimental Results
For training from scratch, each communication round varied the portion of updated layers (75%, 50%, and 25%). Each local client was allocated up to a maximum number of tokens, calculated by dividing the total dataset size by the number of clients, where tokens refer to the units obtained from tokenizing the dataset. A minimum number of tokens is half of this maximum number. Local datasets were non-overlapping, and 10% of clients participated per round. Table I shows that FedTLU consistently achieved better test perplexity, reducing global perplexity by up to 1.14% and local perplexity by up to 1.3% compared to Random. FedTLU also showed a 41.7-120.8% improvement in global perplexity and 13.8-116.2% in local perplexity compared to last-layer updates. These results indicate that targeted layer updates in FedTLU enhance model stability.
For fine-tuning, we trained the global model for 400 rounds with full updates. The same token distribution approach as in training from scratch was applied, ensuring consistency across experiments. During fine-tuning, 5% of clients participated per round, and each client’s local data was halved. The update portion was decreased sequentially from 75% to 50% to 25%, and if the loss did not decrease by at least 1% for 10 consecutive rounds, the portion was further reduced. Table II and Fig. 1 show that FedTLU outperformed other strategies, reducing global perplexity by 2.54-3.50% compared to full updates, by up to 0.97% compared to Random, and by 3.18-3.80% compared to last-layer updates. Locally, FedTLU achieved reductions of 2.60-3.15%, 0.06-0.92%, and 3.23-3.36%. In Fig. 1, FedTLU consistently maintained lower global perplexity during training.